An electronics company makes two types of switches. Type A takes 4 minutes to make and requires ​$3 worth of materials. Type B takes 5 minutes to make and requires ​$5 of materials. In the latest production​batch, it took 32 hours to make these​ switches, and the materials cost ​$1740. How many of each type of switch were​ made?
Answer by ikleyn(53763)

.
4a + 5b = 32*60     (1)     (minutes to make)
3a + 5b = 1740      (2)     (material cost)


To solve the system, use the Elimination method.  For it, subtract eq(2) from eq(1).  You will get


4a - 3a = 32*60 - 1740 = 180.


a =  180.


Then from eq(2),  5b = 1740 - 3*180 =  1200,  b = 1200/5 = 240.


ANSWER.  180 switches of the type A  and   240 switches of the type B.